1 Workpackages

1.1 Workpackage 1: HPF development tools

Leader:  Connexité
Other partners:  Univ. Southampton

BES SA

Duration:  3 years

\underline{Manpower:}

Univ. Southampton 3 MY

task 1.1: A tool for improving ‘dusty deck’ Fortran codes: 1 MY (1 + 0 + 0) task 1.2: Data mapping ‘assistant’: 2 MY (0 + 1 + 1)

BES SA  1.5 MY

task 1.3: Intrinsics and library call generator: 1.5 MY (.5 + .5 + .5)

Connexité  7.5 MY

task 1.4: Adapting Forsys tools for HPF: 1.5 MY (1 + .5 + 0) task 1.5: Building modern interfaces for tools: 3 MY (0 + 1 + 2) task 1.6: Inter-tool protocol: 3 MY (1 + 1.5 + .5)

Total  12 MY

Other resources required:

1.2 Workpackage 2: HPF Visualisation and Graphics

Leader:  Siemens
Other partners:  Univ. Southampton

TUM

Duration:  3 years

\underline{Manpower:}

TUM  3 MY

task 2.1: Trace and cross-reference data: 3 MY (1 + 1 + 1)

Univ. Southampton 3 MY

task 2.2: Array-level visualisation tool: 3 MY (1 + 1 + 1)

Siemens  9 MY

task 2.3: Message-passing and process level visualisation tool: 9 MY (3 + 3 + 3)

IBM Bergen  6 MY

task 2.4: HPF Graphics Library: 6 MY (2 + 2 + 2)

Total  15 MY

Other resources required:

1.3 Workpackage 3: HPF optimisation

Leader:  BES SA
Other partners:  University of Southampton
Duration:  3 years

\underline{Manpower:}

Univ. Southampton 3 MY

task 3.1: HPF translation schemes: 3 MY (1 + 1 + 1)

BES SA  6.5 MY

task 3.2: Source-to-source transformation system: 3 MY (1 + 1 + 1) task 3.3: Source-to-source optimiser: 1.5 MY (0.5 + 0.5 + 0.5) task 3.4: Integrated debugger and profiler: 2 MY (0 + 1 + 1)

Total  9.5 MY

Other resources required:

1.4 Workpackage 4: Evaluation

Leader:  Telmat Informatique
Other partners:  -
Duration:  3 years

\underline{Manpower:}

Telmat  3 MY

task 4.1: Evaluation of project tools: 3 MY (1 + 1 + 1)

Total  3 MY

Other resources required:

1.5 Workpackage 5: Project Management

Leader:  University of Southampton
Other partners:  -
Duration:  3 years

\underline{Manpower:}

Univ. Southampton 1.5 MY

task 5.1: Project Management: 1.5 MY (.5 + .5 + .5)

Total  1.5 MY

Other resources required:

2 Workpackage interdependencies and milestones

4 5 6

3 Major Deliverables and Review Points

3.1 Workpackage 1

Task Deliverable  Id  type availability Months from

project start

1.1: Interprocedural normalisation

Interprocedural normalisation tool  1.1a report  P  12
1.1b tool  P  12

1.2: Data mapping ‘assistant’

Alignment generator  1.2a report  P  24
1.2b tool  P  24
Data mapping ‘assistant’  1.2c report  P  36
1.2d tool  P  36

1.3: Intrinsics / library call generator (ILG)

ILG progress report  1.3a report  ?  12
HPF intrinsics data-base  1.3b tool  ?  12
ILG progress report  1.3c report  ?  24
HPF library data-base  1.3d tool  ?  24
ILG progress report  1.3e report  ?  36
HPF num-library data-base  1.3f tool  ?  36

1.4: Adapting Foresys Tools for HPF

adapting parser, visualizer, Forest  1.4a code  C  6
product Foresys/HPF  1.4b code  C  18

1.5 Building Modern Interfaces for Tools

1.5.1 Source Manipulation Tools

prototype environment w/source tools  1.5a code  R  24
documentation source tools interface  1.5b report  P  36

1.5.2 Run Time Tools 1.5.3 Integration and Packaging

documentation completed HPF environment 1.5c report  P  36
product completed HPF environment  1.5d code  C  36

1.6 Inter-tool Protocol

1.6.1 Defining the communication protocol

protocol specification 1.6a report P 12

1.6.2 Sophtalk implementation of the protocol

product w/existing tools 1.6b code R 18

1.6.3 Chosen standard implementation of the protocol

doc standard version  1.6c report  P  24
product using standard  1.6d code  R  36

3.2 Workpackage 2

Task Deliverable  Id  type availability Months from

project start

2.1: Trace and cross reference data

Definition of trace format, X-ref, interfaces 2.1a report  ?  12
Impln of generation and reading functions 2.1b code  ?  18
Compression of traces and X-refs  2.1c report  ?  24
2.1d code  ?  30
Evaluation  2.1e report  ?  36

2.2: Array-level visualisation tool

Visualisation tool design  2.2a report  P  6
Prototype implementation  2.2b code  P  18
Final implementation  2.2c code  P  30
Evaluation  2.2d code  P  36

2.3: Message-passing and process level visualisation tool

Visualization of PVM/MPI level  2.3a tool  ?  12
HPF execution visualization  2.3b report  P  24
Prototype  2.3c code  ?  24
HPF execution visualization tool  2.3d code  P  30
Evaluation  2.3e report  ?  36

2.4: HPF Graphics Library

Requirements definition document  2.4a report  P  6
First prototype  2.4b code  P  12
User documentation  2.4c report  P  12
Progress report  2.4d report  P  24
Final report  2.4e report  P  36

3.3 Workpackage 3

Task Deliverable  Id  type availability Months from

project start

3.1: Translation Schemes

Literature survey  3.1a report  P  3
Calculation placement strategies  3.1b report  P  12
Progress report  3.1c report  P  24
Testbed implementation and final report tool 3.1d  P  36

3.2: Source-to-source transformation system

progress report  3.2a report  ?  12
batch system  3.2b tool  ?  12
progress report  3.2c report  ?  24
interactive  3.2d tool  ?  24
progress report  3.2e report  ?  36
optimised  3.2f tool  ?  36

3.3: Source-to-source optimiser

progress report  3.3a report  ?  12
prototype data-base  3.3b tool  ?  12
progress report  3.3c report  ?  24
progress report  3.3d report  ?  36
data-base  3.3e tool  ?  36

3.4: Integrated debugger and profiler

progress report  3.4a report  ?  24
prototype  3.4b tool  ?  24
progress report  3.4c report  ?  36
tool  3.4d tool  ?  36

3.4 Workpackage 4

Task Deliverable  Id  type availability Months from

project start

4.1: Ports to HPF

Progress report 4.1a report P 12

4.2: Evaluation of project tools

Final report  4.2a report  P  12

3.5 Workpackage 5

Task Deliverable Id type availability Months from

project start

5.1: Project Management